OpenStudy (anonymous):

Compare these fractions 8/10 12/15

jabez177 (jabez177):

\[15\div5\]

OpenStudy (gabbyalicorn):

Do you think its greater than, equal or less than hunter?

OpenStudy (gabbyalicorn):

Nvm, he's offline.

jabez177 (jabez177):

\[15\div5\]

jabez177 (jabez177):

12 divided by 2 = ? 8 divided by 2 =?

OpenStudy (anonymous):

I was talking to me teacher

jabez177 (jabez177):

\[10\div5=2\]\[15\div5=3\]

jabez177 (jabez177):

\[8\div2=4\]\[12\div2=6\]

jabez177 (jabez177):

\[\frac{ 4 }{ 6 }\] can be simplified further. 4 divided by 2 is 2 6 dividde by 2 is 3 So 2/3

jabez177 (jabez177):

Is \[\frac{ 2 }{ 3 }=\frac{ 2 }{ 3 }\] ?